Calculating Poisson Probabilities (Edexcel International A Level (IAL) Maths: Statistics 2): Revision Note

Exam code: YMA01

Amber

Written by: Amber

Reviewed by: Dan Finlay

Updated on

Guided study available on this topic

Understand this topic with deeper explanations and understanding checks.

Calculating Poisson Probabilities

Throughout this section we will use the random variable X~Po(λ)  . For a Poisson distribution, the probability of a X taking a non-integer or negative value is always zero. Therefore any values mentioned in this section (other than λ) will be assumed to be non-negative integers.

Where does the formula for a Poisson distribution come from?

  • The formula for calculating an individual Poisson probability is

    • P(X=r)=pr=e−λ×λrr! for r = 0, 1, 2 ,...., n

  • The derivation of the formula comes from the binomial distribution, however it is outside the scope of this syllabus and will not be proved here

    • Whilst the binomial distribution relies on knowing a fixed number of trials, the Poisson can allow for any number of trials within a time period

    • Only the mean number of occurrences of an event within the given period needs to be known

How do I calculate the cumulative probabilities for a Poisson distribution? 

  • Most of the time the value of λ will be in the table for the 'Poisson Cumulative Distribution Function' in the formula booklet

  • P(X≤x) is asking you to find the probabilities of all values up to and including x   

    • This will be the value in the row of x

  • P(X<x) is asking you to find the probabilities of all values up to but not including x

    • This means all values that are less than x

    • This will be the value in the row of x - 1

      • P(X≤x−1)

  • P(X≥x) is asking you to find the probabilities of all values greater than and including x

    • This means all values that are at least x

    • As there is not a fixed number of trials this includes an infinite number of possibilities

    • To calculate this, use the identity P(X≥x)=1−P(X<x)

      • Using above we get: P(X≥x)=1−P(X≤x−1)

  • P(X>x) is asking you to find the probabilities of all values greater than but not including x 

    • This means all values that are more than x

    • Rewrite P(X>x)=1−P(X≤x) as to calculate this

  • If calculating P(a≤X≤b) then calculate P(X≤b)−P(X≤a−1)

    • This is the same idea as for the binomial distribution

How do I calculate the cumulative probabilities for a Poisson distribution if λ is not in the table?

  • In the unlikely case that λ is not in the table

    • Use the formula to find the individual probabilities and then add them up

    • Make sure you are confident working with inequalities for discrete values

    • Only integer values will be included so it is easiest to look at which integer values you should include within your calculation

    • Sometimes it is quicker to find the probabilities that are not being asked for and subtract from one

  • Having to type a lot of calculations into your calculator can be time consuming and cause errors

  • Consider the calculation P(X≤3)=e−λ×λ00!+e−λ×λ11!+e−λ×λ22!+e−λ×λ33!

    • Note that e−λ exists in every term and can be factorised out

    • Recall that λ0=1 and 0! = 1

    • Recall also that λ1=λ and 1! = 1

  • This calculation could be factorised and simplified to

    • P(X≤3)=e−λ(1+λ+λ22!+λ33!) 

    • This is much simpler and easier to type into your calculator in exam conditions

How do I change the mean for a Poisson distribution?

  • Sometimes the mean may be given for a different interval of time or space than that which you need to calculate the probability for

  • A given value of λ can be adjusted to fit the necessary time period

    • For example if a football team score a mean of 2 goals an hour and we want to find the probability of them scoring a certain number of goals in a 90 minute game, then we would use the distribution X ~ Po(3)   

      • 90 = 1.5 × 60 so use 1.5λ

Worked Example

Xiao makes silly mistakes in his maths homework at a mean rate of 2 per page.

(a) Define a suitable distribution to model the number of silly mistakes Xiao would make in a piece of homework that is five pages long.

(b) Find the probability that in any random page of Xiao’s homework book there are

Answer:

(i) exactly three silly mistakes

(ii) at most three silly mistakes

(iii) more than three silly mistakes.

1-2-2-calculating-poisson-probabilities-we-solution-part-1
1-2-2-calculating-poisson-probabilities-we-solution-part-2

Examiner Tips and Tricks

  • Look carefully at the given time or space interval to check if you need to change the mean before carrying out calculations.

    • Be prepared for this to change between question parts!

Unlock more, it's free!

Join the 100,000+ Students that ❤️ Save My Exams

the (exam) results speak for themselves:

Build on this topic

Amber

Author: Amber

Expertise: Maths Content Creator

Amber gained a first class degree in Mathematics & Meteorology from the University of Reading before training to become a teacher. She is passionate about teaching, having spent 8 years teaching GCSE and A Level Mathematics both in the UK and internationally. Amber loves creating bright and informative resources to help students reach their potential.

Dan Finlay

Reviewer: Dan Finlay

Expertise: Portfolio Lead

Dan graduated from the University of Oxford with a First class degree in mathematics. As well as teaching maths for over 8 years, Dan has marked a range of exams for Edexcel, tutored students and taught A Level Accounting. Dan has a keen interest in statistics and probability and their real-life applications.